3. Only 4% of clover plants are of the four-leaf variety. You will pick clover plants at random until you find a four-leaf clover.

media%2Fd75%2Fd754d5fb-eeaf-403b-b272-ac
Answer #1

a)

p = 0.04

X = number of plants picked to get first 4-leaf clover

X follows geometric distribution

P(X = k) = (1-p)^(k-1)* p

P(X =16) = 0.96^15 * 0.04

= 0.02168

__________________________________

b)

mean = 1/p = 1/0.04 = 25

sd = sqrt((1-p)/p^2) = sqrt(0.96/0.04^2)

= 24.494897